Let S \subset \mathbb{R}^k be a set. Show that the convex hull of S, denoted by cvx(S), is equal to the set \bigcup \left\{ (1-t)x + tx \, | \, t \in [0,1], x,y \in S \right\}



cvx(S)
Answer #1

Convex hull of a set S is the set of all convex linear combinations of elements in S. Let x,y \in S ; t \in [0,1] then (1-t)x+ty \in \cup_t \left \{ (1-t)x+ty| t \in [0,1],x,y \in S \right \} which is a convex linear combination of x and y hence it is an element in cvx(S). Hence every element of the given union is an element of the convex hull. Therefore \cup_t \left \{ (1-t)x+ty| t \in [0,1],x,y \in S \right \} \subseteq cvx(S) ----(1)

Now let x is an element in S then x = 1.x | 1 \in [0,1] \Rightarrow x \in \cup_t \left \{ (1-t)x+ty| t \in [0,1],x,y \in S \right \} , thereforeS\subseteq \cup_t \left \{ (1-t)x+ty| t \in [0,1],x,y \in S \right \}. Now the union contains all possible convex linear combination of elements in S and cvx(S) is the smallest set containing all such elements thus we have cvx(S)\subseteq \cup_t \left \{ (1-t)x+ty| t \in [0,1],x,y \in S \right \} ----(2).

From (1) and (2) we get cvx(S)=\cup_t \left \{ (1-t)x+ty| t \in [0,1],x,y \in S \right \} .